OpenStudy (bethgrace96):

medal! i will fan u. The point 3,4 lies on the graph of f(x)= sqrt(2x+a). Which ordered pair below lies on the graph of g(x)=1/2sqrt(2x+a)? a)3/2,2 b)5/2,3/2 c)3,2 d)27,8

zepdrix (zepdrix):

So you need to figure out what a is,\[\Large\rm \color{royalblue}{f(x)}=\sqrt{2\color{orangered}{x}+a}\]And they gave us the ordered pair,\[\Large\rm \left(\color{orangered}{3},~\color{royalblue}{4}\right)\]So when you plug those values in, do you understand how to solve for a?\[\Large\rm \color{royalblue}{4}=\sqrt{2\color{orangered}{(3)}+a}\]

OpenStudy (bethgrace96):

so whats the answer

OpenStudy (bethgrace96):

?????????

OpenStudy (anonymous):

okay so using @zepdrix 's formula that he has given you, you would solve like this: